# Identifiable learning of dissipative dynamics

### v1

@@ Appendix A Theoretical Results
In this section, we present the theoretical results concerning the I-OnsagerNets. We consider an Itô SDE driven by standard Brownian motion
@@
where $g$ and $\sigma$ denote the drift and diffusion coefficients, respectively. To ensure the well-posedness of (9) and the classical validity of its stationary Fokker-Planck equation, we adopt the following standard technical assumptions:
@@ 1. Linear growth: There exists a constant $K_{g}>0$ such that for all $\mathbf{Z}\in\mathbb{R}^{D}$,
Under these conditions, (9) admits a unique strong solution [64].
@@
For convenience, we repeat the dynamics of I-OnsagerNet:
@@ A.1 Stationary Density
solves the stationary Fokker–Planck equation:
@@
A.2 Time-reversal and Entropy Production
The theory of time-reversed diffusions was originally developed by Hausmann and Pardoux [34]. It was subsequently extended by Millet, Nualart, and Sanz under modified conditions [56]. We adopt the framework of Da Costa and Pavliotis [19], which applies to SDEs with locally Lipschitz coefficients.
@@ A.2 Time-reversal and Entropy Production
With the above lemma, we now able to derive the time-reversed diffusion process for the I-OnsagerNet. Substituting (12), (13), and (15), we obtain:
@@
A.2.1 Global EPR
By the time-reversed diffusions, we can calculate the global EPR. A detailed derivation is available in [39], and its extension to cases with a noninvertible diffusion matrix is treated in [19]. To make this SI self-contained, we briefly recount the derivation steps.
@@ A.2.1 Global EPR
A.2.2 Current Velocity and System EPR
The time-irreversible drift plays a fundamental role in the dynamics of the system. Besides being used in computing the global EPR, $f_{\text{irr}}$ is identical to the current velocity [58, 61, 65], as shown by the equivalence:
@@ A.2.2 Current Velocity and System EPR
defines the characteristics associated with the stationary FPE given by (14). The resulting flow map $\phi_{t}$, defined via $\phi_{t}(\mathbf{Z})=\mathbf{Z}_{t}^{\text{irr}}$ where $\mathbf{Z}_{t}^{\text{irr}}$ solves (16) with initial condition $\mathbf{Z}_{0}^{\text{irr}}=\mathbf{Z}$, preserves the stationary density. This implies that for any observable $A(\mathbf{Z})$, the following identity holds:
@@
We then show that the system EPR can be computed from $f_{\text{irr}}$. The system EPR quantifies the trajectory-wise evolution of the stochastic Gibbs entropy, defined as [65, 66]:
@@
Substituting the stationary density $\rho$ from (13), we obtain:
@@
where $\nabla\cdot(\nabla\cdot W)=0$ because the second divergence of a anti-symmetric matrix field vanishes. In summary, for I-OnsagerNet, the system EPR has several equivalent expressions:
@@
A.3 Mathematical Characterization
In this section, we prove the expressivity and uniqueness of I-OnsagerNet. Our arguments mainly rely on the Helmholtz decomposition [19, 25, 31, 39, 61] and the essentially Hamiltonian decomposition [26]. The main theorem is repeated below. Recall that $M=\sigma\sigma^{\top}/2$.
@@ A.3 Mathematical Characterization
For any SDE as given by (9), there exists a banded anti-symmetric matrix $W$ with bandwidth 1 and a potential $V(\mathbf{Z})$ satisfying $e^{-V}\in L^{1}(\mathbb{R}^{D})$ such that the drift $g$ admits the following decomposition:
@@
This result, known as the essentially Hamiltonian decomposition, was proved by Feng and Shang [26] for constructing volume-preserving integrators for divergence-free ODEs. For completeness, we provide its proof.
@@
where $M$ is the diffusion matrix and $\rho$ is the stationary density. The time-irreversible part is then given by $g_{\text{irr}}=g-g_{\text{rev}}$. Since $\rho$ satisfies the stationary Fokker–Planck equation [61]:
@@
Next we prove the uniqueness. We have shown in Section A.1 that the density
@@
is a stationary density of the system. By the uniqueness assumption of the stationary density, the potential $V$ in the (10) is uniquely determined up to an additive constant and satisfies
@@
Appendix B Properties of Global EPR
In this section, we briefly review the necessary properties of global EPR.
B.1 Coarse-graining by Projection
@@
This decomposition is a well-established result in information theory [16]. For completeness, we provide a proof in the case where densities exist.
@@
B.2 Direct Numerical Computation
In this section, we present a numerical method for computing the global EPR.
@@ Appendix C Barrier Heights
The barrier height for the transition from a stable state $A$ to another stable state $B$ is defined as the minimal additional value of an effective energy function $E$ that the system must acquire to escape from $A$ along the optimal transition path connecting $A$ and $B$ [33, 71]. Formally, let $E^{\ddagger}$ denote the maximum energy along the minimum energy path between $A$ and $B$; the barrier height is then given by
@@
To numerically compute the barrier height between two stable states, we proceed as follows. First, we select two points located near the respective stable states. Each given point is first relaxed by gradient descent to determine its corresponding local minimum, identified as the stable states of interest. Then we seek the first-order saddle point on the minimum-energy path connecting the two minima. This is achieved using the Gentlest Ascent Dynamics (GAD) method [74], which modifies the gradient flow so that the dynamics climb along the unstable direction while descending along all others. Specifically, for a system with potential energy $E(x)$, GAD evolves both the position $x$ and a direction vector $v$ according to
@@
Typically, the effective energy function is characterized by the quasi-potential. In the small noise limit, it can be asymptotically approximated by the negative logarithm of the stationary density, scaled by the noise intensity [5, 27]. In this paper, I-OnsagerNets yield the negative logarithm of the stationary density under varying external forces at fixed temperature, ensuring that the resulting trends in barrier heights are consistent with the theoretical definition. Although the energy function returned by OnsagerNet lacks a strict physical definition, we estimate its barrier height using the same procedure for comparative analysis.
Appendix D Supplementary Numerical Results
D.1 Linear Case
@@
D.2 Polymer Stretching Dynamics
D.2.1 Simulation Data Preparation
This section describes the workflow for generating the simulation dataset, including modelling and simulation of polymer dynamics at the microscopic scale and computing representative coordinates at the macroscopic scale.
Microscopic polymer dynamics simulation. We begin by describing the model for microscopic polymer dynamics. The essential features of the governing equations are summarized below; additional details can be found in [13].
@@ D.2.1 Simulation Data Preparation
where $r_{ij}$ denotes the distance between beads $i$ and $j$, and $\mu=35\ \mathrm{pN}$ is chosen to effectively suppress chain crossings, as validated in [73]. Constraint forces enforce fixed bond lengths and are expressed as:
@@
where $\mathbf{b}_{i}$ is the unit vector along bond $i$, and $T_{i}$ is the tension in that bond. Brownian forces $\mathbf{F}^{\mathrm{br}}_{i}(t)$ are modelled as Gaussian white noise satisfying the fluctuation–dissipation theorem:
@@
Numerical integration is performed using a predictor–corrector algorithm [52]. The rigid-rod constraints lead to a non-linear system of equations for the tensions $T_{i}$, which is solved at each time step via Newton’s method [69].
@@
We outline the procedure for obtaining the remaining closure coordinates within a dimensionality reduction framework. Specifically, we apply the PCA–ResNet approach [13] to extract these coordinates:
@@
Although the training data are derived from explicit bead–rod simulations, I-OnsagerNet operates only on reduced coordinates. This property decouples the learning framework from system-specific details, allowing its application to a broad range of non-equilibrium systems.
D.2.2 Experimental Data Preparation
@@
An automated single-molecule stretching trap is implemented to obtain experimental image data, as originally developed in [68]. In the following, we introduce the key aspects of the setup. An illustration of the experimental data preparation is provided in Fig. 6. Comprehensive descriptions of the experimental protocol and material preparation are available in [68].
@@
To enable long-term observation and controlled stretching, an active feedback system is implemented to stabilize the saddle point. Images are acquired at 50 ms intervals using an inverted fluorescence microscope equipped with a high-resolution sCMOS camera. A custom LabVIEW program processes the image stream in real time to extract the molecular centroid and projected contour length. A proportional feedback loop with a gain of $G=2.2~\mathrm{V}/\mu\mathrm{m}$ adjusted the East–West voltage to maintain the DNA centroid within $1~\mu\mathrm{m}$ of the saddle point during stretching, thereby allowing stable trapping and extension of the molecule in a planar elongational field over extended periods.
@@
Figure 8: The total EPR $\dot{s}_{\text{tot}}$ on test trajectories projected onto the $Z_{1}\text{–}Z_{2}$ and $Z_{1}\text{–}Z_{3}$ planes. The background is the corresponding potential energy surface. One example trajectory from the test dataset is overlaid on the $Z_{1}\text{–}Z_{3}$ panels.
D.2.3 Additional Numerical Results
@@
To assess the accuracy of the learned I-OnsagerNet, we evaluate its performance on both the training and test trajectories. Moreover, validation is performed using two distinct initial chain conformations absent from the training set. These representative configurations were deliberately initialized with near-identical projected extensions yet diverged significantly in their transient stretching dynamics, with one stretching much faster than the other. As shown in Fig. 7, on both the training and test datasets across all cases, I-OnsagerNet precisely captures these contrasting behaviors and reproduces the true ensemble evolution of chain extension under varying external force. I-OnsagerNet also achieves high-fidelity reconstruction of the first-passage time distributions required to attain a specified extension threshold.
@@
D.3 Stochastic Gradient Langevin Dynamics
D.3.1 Least Squares Regression
In this paper, we leverage I-OnsagerNet to learn the dynamics of the SGLD discrete-time process. As an example, consider the least squares problem:
@@ D.3.1 Least Squares Regression
where $\{\gamma_{1},\gamma_{2},\ldots,\gamma_{b}\}$ is a randomly selected set of indices. $\boldsymbol{\xi}^{(t)}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is standard Gaussian noise. It is remarked that when $n=1$ and noise is omitted, this reduces to the randomized Kaczmarz method [40, 70]. And we use it as a test case for evaluating linear SGLD dynamics.
@@
D.3.2 Independent Components Analysis
In the following, we provide an overview of a widely used independent component analysis (ICA) algorithm based on stochastic natural-gradient optimization [1], and present its SGLD version [75]. ICA aims to recover statistically independent latent sources from their linear mixtures. It assumes that the observed signal $\mathbf{x}\in\mathbb{R}^{D}$ is generated as $\mathbf{x}=W^{-1}\mathbf{s}$, where the latent components $\mathbf{s}=(s_{1},\dots,s_{D})$ are mutually independent and typically exhibit non-Gaussian, heavy-tailed marginal distributions. A probabilistic formulation of ICA thus introduces the likelihood
@@ D.3.2 Independent Components Analysis
Following [1], optimization of $W$ employs natural-gradient descent, which preconditions the Euclidean gradient by $W^{\mathsf{T}}W$ to account for the Riemannian geometry of the parameter manifold. In particular, if we choose $p_{d}(y_{d})=1/(4\cosh^{2}(y_{d}/2))$ with $y_{d}=\mathbf{w}_{d}^{\mathsf{T}}\mathbf{x}$, the gradient takes the form:
@@
where $\boldsymbol{\xi}^{(t)}\in\mathbb{R}^{D\times D}$ has i.i.d. entries drawn from $\mathcal{N}(0,1)$ and $\eta$ denotes the step size. The multiplicative factor $\sqrt{W_{t}^{\mathsf{T}}W_{t}}$ reflects the natural-gradient metric preconditioning. In the SGLD [75] formulation, the full-data gradient is replaced by a mini-batch approximation of size $b$, yielding the update rule:
@@
D.3.3 Additional Numerical Results
@@
To evaluate the robustness and reconstruction fidelity of I-OnsagerNet in the context of least-squares regression with SGLD, we examine its ability to reproduce both transient trajectories and output losses across diverse parameter settings (Fig. 9). I-OnsagerNet consistently recovers the temporal evolution of the stochastic dynamics under different batch sizes and initialization schemes, demonstrating that the learned representation captures the underlying generative mechanism rather than overfitting to specific training conditions. This robustness enables a mechanistic interpretation of SGLD behavior directly from the learned model.
We further validate I-OnsagerNet on the nonlinear ICA benchmark (Fig. 10). Across varying batch sizes and initialization schemes, the model yields highly consistent reconstructions of both the transient trajectories and the invariant distributions. These results confirm that I-OnsagerNet accurately captures the stochastic dynamics of the ICA with SGLD.
D.3.4 Theoretical Justification for EPR as a Measure of Sampling Error
For convenience, we recall that the discrete-time dynamics of SGLD converge weakly, in the vanishing step-size limit, to a continuous-time diffusion process governed by the Itô SDE [48]:
@@ D.3.4 Theoretical Justification for EPR as a Measure of Sampling Error
where $\{\gamma_{i}\}_{i=1}^{b}$ indexes a mini-batch of size $b$. By Theorem 1, there exists a matrix field $W_{b}(\mathbf{Z})$ of the form (11) and a potential function $V_{b}(\mathbf{Z})$ such that the drift term can be decomposed as:
@@
RFI provides an upper bound for the Wasserstein distance and the KL divergence according to HWI inequality [60] under the Bakry-Émery condition, with applications in information geometry and statistical physics [72, 57]. For SGLD, the RFI is given by
@@
The difference between (18) and (19) is attributed to the $\Sigma_{b}$ term, reflecting the effect of minibatch-induced stochasticity and diminishing as the batch size increases. This observation provides a theoretical justification for the empirical finding that both the sampling error and the global EPR consistently decrease with larger batch sizes. In addition, since all quantities involved ($V_{b}$, $W_{b}$, $M_{b}$, and consequently $f_{\mathrm{irr}}$ and $\Sigma_{b}$) are explicitly parameterized and estimable within our framework I-OnsagerNet, (19) can be computed directly from data and thus serves as a principled scalar indicator of sampling error in SGLD.

### v2

@@ Appendix B Theoretical Results
In this subsection, we present the theoretical results concerning the UIDDs. We consider an Itô SDE driven by standard Brownian motion
@@
where $g$ and $\sigma$ denote the drift and diffusion coefficients, respectively. To ensure the well-posedness of Eq. (8) and the classical validity of its stationary Fokker-Planck equation, we adopt the following standard technical assumptions:
@@ 1. Linear growth: There exists a constant $K_{g}>0$ such that for all $\mathbf{Z}\in\mathbb{R}^{D}$,
Under these conditions, Eq. (8) admits a unique strong solution [prevot2007concise].
@@
For convenience, we repeat the dynamics of UIDD:
@@ B.1 Stationary Density
solves the stationary Fokker-Planck equation:
@@
B.2 Time-reversal and Entropy Production
The theory of time-reversed diffusions was originally developed by Hausmann and Pardoux [haussmann1986time]. It was subsequently extended by Millet, Nualart, and Sanz under modified conditions [millet1989integration]. We adopt the framework of Da Costa and Pavliotis [da2023entropy], which applies to SDEs with locally Lipschitz coefficients.
@@ B.2 Time-reversal and Entropy Production
With the above lemma, we now able to derive the time-reversed diffusion process for the UIDD. Substituting Eq. (11), Eq. (12), and Eq. (14), we obtain:
@@
Global EPR
By the time-reversed diffusions, we can calculate the global EPR. A detailed derivation is available in [jiang2004mathematical], and its extension to cases with a noninvertible diffusion matrix is treated in [da2023entropy]. To make this SI self-contained, we briefly recount the derivation steps.
@@
Current Velocity and System EPR
The time-irreversible drift plays a fundamental role in the dynamics of the system. Besides being used in computing the global EPR, $f_{\text{irr}}$ is identical to the current velocity [nelson1967dynamical, pavliotis2014stochastic, seifert2005entropy], as shown by the equivalence:
@@
defines the characteristics associated with the stationary FPE given by Eq. (13). The resulting flow map $\phi_{t}$, defined via $\phi_{t}(\mathbf{Z})=\mathbf{Z}_{t}^{\text{irr}}$ where $\mathbf{Z}_{t}^{\text{irr}}$ solves Eq. (15) with initial condition $\mathbf{Z}_{0}^{\text{irr}}=\mathbf{Z}$, preserves the stationary density. This implies that for any observable $A(\mathbf{Z})$, the following identity holds:
@@
We then show that the system EPR can be computed from $f_{\text{irr}}$. The system EPR quantifies the trajectory-wise evolution of the stochastic Gibbs entropy, defined as [seifert2005entropy, seifert2012stochastic]:
@@
Substituting the stationary density $\rho$ from Eq. (12), we obtain:
@@
where $\nabla\cdot(\nabla\cdot W)=0$ because the second divergence of a anti-symmetric matrix field vanishes. In summary, for UIDD, the system EPR has several equivalent expressions:
@@
B.3 Mathematical Characterization
In this subsection, we prove the expressivity and uniqueness of UIDD. Our arguments mainly rely on the Helmholtz decomposition [da2023entropy, eyink1996hydrodynamics, graham1977covariant, jiang2004mathematical, pavliotis2014stochastic] and the essentially Hamiltonian decomposition [feng1995volume]. The main theorem is repeated below. Recall that $M=\sigma\sigma^{\top}/2$.
@@ B.3 Mathematical Characterization
For any SDE as given by Eq. (8), there exists a banded anti-symmetric matrix $W$ with bandwidth 1 and a potential $V(\mathbf{Z})$ satisfying $e^{-V}\in L^{1}(\mathbb{R}^{D})$ such that the drift $g$ admits the following decomposition:
@@
This result, known as the essentially Hamiltonian decomposition, was proved by Feng and Shang [feng1995volume] for constructing volume-preserving integrators for divergence-free ODEs. For completeness, we provide its proof.
@@
where $M$ is the diffusion matrix and $\rho$ is the stationary density. The time-irreversible part is then given by $g_{\text{irr}}=g-g_{\text{rev}}$. Since $\rho$ satisfies the stationary Fokker-Planck equation [pavliotis2014stochastic]:
@@
Next we prove the uniqueness. We have shown in subsection B.1 that the density
@@
is a stationary density of the system. By the uniqueness assumption of the stationary density, the potential $V$ in the Eq. (9) is uniquely determined up to an additive constant and satisfies
@@
Appendix C Algorithm
In this subsection, we detail the architecture of the UIDD modules and outline the practical implementation for learning.
C.1 Neural Network Parameterization
$W(\mathbf{Z})=\sum_{d=1}^{D-1}H_{d}(\mathbf{Z})J_{d},$
where $U:\mathbb{R}^{D}\rightarrow\mathbb{R}^{m}$ is a neural network, $\gamma_{ij}$ are trainable parameters, and $\beta_{1},\beta_{2},\beta_{3}>0$ are fixed scaling factors. This architecture ensures that the potential function satisfies both the integrability condition and the appropriate growth requirements specified in Methods B. The expressivity is achieved through the embedded neural network $U$ within this constrained functional form. In our implementation, we set $m=32$ and use a network $U$ with one hidden layer of 128 neurons. We employ the shifted ReQU activation in polymer dynamics and tanh in the SGLD context.
C.2 Learning Algorithm
Each trajectory is assumed to be a realization of an underlying SDE starting from the initial condition $\mathbf{Z}^{(n)}_{0}$. Our objective is to learn the functions $V$, $W$, and $\sigma$ that govern the dynamics of the UIDD within this deep learning framework.
As described in the previous subsection, we model $V_{\theta}$, $W_{\theta}$, and $\sigma_{\theta}$ with neural networks, resulting in the following parametric SDE:
where
We will use the Maximum Likelihood Estimation (MLE) to learn the parameter $\theta$ [pavliotis2014stochastic, Sec. 5.3]. Let $p(\Delta t,\mathbf{Z}\mid 0,\mathbf{Z}_{0};\theta)$ be the transition density of the process (i.e., the conditional probability density of the system being in state $\mathbf{Z}$ at time $\Delta t$, given it started at $\mathbf{Z}_{0}$ at time 0). The MLE objective is:
where $\Delta t_{m}=t_{m}-t_{m-1}$. Since the exact transition density is typically intractable for general nonlinear diffusion processes, we approximate it using a numerical approximation $p_{h}(\Delta t,\mathbf{Z}\mid 0,\mathbf{Z}_{0};\theta)$ based on the method developed in [zhu2024dyngma]. The practical optimization problem thus becomes:
Specifically, the one-step scheme from [zhu2024dyngma] leads to a Gaussian approximation for the transition density:
$p_{h}(\Delta t,\mathbf{Z}\mid 0,\mathbf{Z}_{0};\theta)$
This approximation, corresponding to the Euler-Maruyama scheme, has been widely used in related works [chen2024constructing, dietrich2023learning]. We employ the Adam optimizer [kingma2014adam] for training, with detailed hyperparameter settings provided in Methods D.
Appendix D Supplementary Information for Numerical Results
D.1 Properties of Global EPR
In this subsection, we briefly review the necessary properties of global EPR.
Coarse-graining by Projection
@@ D.1 Properties of Global EPR
This decomposition is a well-established result in information theory [cover1999elements]. For completeness, we provide a proof in the case where densities exist.
@@
Direct Numerical Computation
In this subsection, we present a numerical method for computing the global EPR.
@@ D.2 Barrier Heights
The barrier height for the transition from a stable state $A$ to another stable state $B$ is defined as the minimal additional value of an effective energy function $E$ that the system must acquire to escape from $A$ along the optimal transition path connecting $A$ and $B$ [hanggi1990reaction, van1992stochastic]. Formally, let $E^{\ddagger}$ denote the maximum energy along the minimum energy path between $A$ and $B$; the barrier height is then given by
@@
To numerically compute the barrier height between two stable states, we proceed as follows. First, we select two points located near the respective stable states. Each given point is first relaxed by gradient descent to determine its corresponding local minimum, identified as the stable states of interest. Then we seek the first-order saddle point on the minimum-energy path connecting the two minima. This is achieved using the Gentlest Ascent Dynamics (GAD) method [weinan2011gentlest], which modifies the gradient flow so that the dynamics climb along the unstable direction while descending along all others. Specifically, for a system with potential energy $E(x)$, GAD evolves both the position $x$ and a direction vector $v$ according to
@@
Typically, the effective energy function is characterized by the quasi-potential. In the small noise limit, it can be asymptotically approximated by the negative logarithm of the stationary density, scaled by the noise intensity [biswas2009small, freidlin2012random]. In this paper, UIDDs yield the negative logarithm of the stationary density under varying external forces at fixed temperature, ensuring that the resulting trends in barrier heights are consistent with the theoretical definition. Although the energy function returned by OnsagerNet lacks a strict physical definition, we estimate its barrier height using the same procedure for comparative analysis.
D.3 Linear Case
@@
D.4 Polymer Stretching Dynamics
Simulation Data Preparation
This subsection describes the workflow for generating the simulation dataset, including modelling and simulation of polymer dynamics at the microscopic scale and computing representative coordinates at the macroscopic scale.
Microscopic polymer dynamics simulation. We begin by describing the model for microscopic polymer dynamics. The essential features of the governing equations are summarized below; additional details can be found in [chen2024constructing].
@@ D.4 Polymer Stretching Dynamics
where $r_{ij}$ denotes the distance between beads $i$ and $j$, and $\mu=35\ \mathrm{pN}$ is chosen to effectively suppress chain crossings, as validated in [vologodskii2006brownian]. Constraint forces enforce fixed bond lengths and are expressed as:
@@
where $\mathbf{b}_{i}$ is the unit vector along bond $i$, and $T_{i}$ is the tension in that bond. Brownian forces $\mathbf{F}^{\mathrm{br}}_{i}(t)$ are modelled as Gaussian white noise satisfying the fluctuation-dissipation theorem:
@@
Numerical integration is performed using a predictor-corrector algorithm [liu1989flexible]. The rigid-rod constraints lead to a non-linear system of equations for the tensions $T_{i}$, which is solved at each time step via Newton’s method [somasi2002brownian].
@@
We outline the procedure for obtaining the remaining closure coordinates within a dimensionality reduction framework. Specifically, we apply the PCA-ResNet approach [chen2024constructing] to extract these coordinates:
@@
Although the training data are derived from explicit bead-rod simulations, UIDD operates only on reduced coordinates. This property decouples the learning framework from system-specific details, allowing its application to a broad range of non-equilibrium systems.
Experimental Data Preparation
@@
An automated single-molecule stretching trap is implemented to obtain experimental image data, as originally developed in [soh2023automated]. In the following, we introduce the key aspects of the setup. An illustration of the experimental data preparation is provided in Fig. 6. Comprehensive descriptions of the experimental protocol and material preparation are available in [soh2023automated].
@@
To enable long-term observation and controlled stretching, an active feedback system is implemented to stabilize the saddle point. Images are acquired at 50 ms intervals using an inverted fluorescence microscope equipped with a high-resolution sCMOS camera. A custom LabVIEW program processes the image stream in real time to extract the molecular centroid and projected contour length. A proportional feedback loop with a gain of $G=2.2~\mathrm{V}/\mu\mathrm{m}$ adjusted the East-West voltage to maintain the DNA centroid within $1~\mu\mathrm{m}$ of the saddle point during stretching, thereby allowing stable trapping and extension of the molecule in a planar elongational field over extended periods.
@@
Figure 8: The total EPR $\dot{s}_{\text{tot}}$ on test trajectories projected onto the $Z_{1}\text{-}Z_{2}$ and $Z_{1}\text{-}Z_{3}$ planes. The background is the corresponding potential energy surface. One example trajectory from the test dataset is overlaid on the $Z_{1}\text{-}Z_{3}$ panels.
Additional Numerical Results
@@
To assess the accuracy of the learned UIDD, we evaluate its performance on both the training and test trajectories. Moreover, validation is performed using two distinct initial chain conformations absent from the training set. These representative configurations were deliberately initialized with near-identical projected extensions yet diverged significantly in their transient stretching dynamics, with one stretching much faster than the other. As shown in Fig. 7, on both the training and test datasets across all cases, UIDD precisely captures these contrasting behaviors and reproduces the true ensemble evolution of chain extension under varying external force. UIDD also achieves high-fidelity reconstruction of the first-passage time distributions required to attain a specified extension threshold.
@@
D.5 Stochastic Gradient Langevin Dynamics
Least Squares Regression
In this paper, we leverage UIDD to learn the dynamics of the SGLD discrete-time process. As an example, consider the least squares problem:
@@ D.5 Stochastic Gradient Langevin Dynamics
where $\{\gamma_{1},\gamma_{2},\ldots,\gamma_{b}\}$ is a randomly selected set of indices. $\boldsymbol{\xi}^{(t)}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is standard Gaussian noise. It is remarked that when $n=1$ and noise is omitted, this reduces to the randomized Kaczmarz method [karczmarz1937angenaherte, strohmer2009randomized]. And we use it as a test case for evaluating linear SGLD dynamics.
@@
Independent Components Analysis
In the following, we provide an overview of a widely used independent component analysis (ICA) algorithm based on stochastic natural-gradient optimization [amari1995new], and present its SGLD version [welling2011bayesian]. ICA aims to recover statistically independent latent sources from their linear mixtures. It assumes that the observed signal $\mathbf{x}\in\mathbb{R}^{D}$ is generated as $\mathbf{x}=W^{-1}\mathbf{s}$, where the latent components $\mathbf{s}=(s_{1},\dots,s_{D})$ are mutually independent and typically exhibit non-Gaussian, heavy-tailed marginal distributions. A probabilistic formulation of ICA thus introduces the likelihood
@@
Following [amari1995new], optimization of $W$ employs natural-gradient descent, which preconditions the Euclidean gradient by $W^{\mathsf{T}}W$ to account for the Riemannian geometry of the parameter manifold. In particular, if we choose $p_{d}(y_{d})=1/(4\cosh^{2}(y_{d}/2))$ with $y_{d}=\mathbf{w}_{d}^{\mathsf{T}}\mathbf{x}$, the gradient takes the form:
@@
where $\boldsymbol{\xi}^{(t)}\in\mathbb{R}^{D\times D}$ has i.i.d. entries drawn from $\mathcal{N}(0,1)$ and $\eta$ denotes the step size. The multiplicative factor $\sqrt{W_{t}^{\mathsf{T}}W_{t}}$ reflects the natural-gradient metric preconditioning. In the SGLD [welling2011bayesian] formulation, the full-data gradient is replaced by a mini-batch approximation of size $b$, yielding the update rule:
@@
Additional Numerical Results
@@
To evaluate the robustness and reconstruction fidelity of UIDD in the context of least-squares regression with SGLD, we examine its ability to reproduce both transient trajectories and output losses across diverse parameter settings (Fig. 9). UIDD consistently recovers the temporal evolution of the stochastic dynamics under different batch sizes and initialization schemes, demonstrating that the learned representation captures the underlying generative mechanism rather than overfitting to specific training conditions. This robustness enables a mechanistic interpretation of SGLD behavior directly from the learned model.
We further validate UIDD on the nonlinear ICA benchmark (Fig. 10). Across varying batch sizes and initialization schemes, the model yields highly consistent reconstructions of both the transient trajectories and the invariant distributions. These results confirm that UIDD accurately captures the stochastic dynamics of the ICA with SGLD.
Theoretical Justification for EPR as a Measure of Sampling Error
For convenience, we recall that the discrete-time dynamics of SGLD converge weakly, in the vanishing step-size limit, to a continuous-time diffusion process governed by the Itô SDE [li2019stochastic]:
@@
where $\{\gamma_{i}\}_{i=1}^{b}$ indexes a mini-batch of size $b$. By Theorem 1, there exists a matrix field $W_{b}(\mathbf{Z})$ of the form Eq. (10) and a potential function $V_{b}(\mathbf{Z})$ such that the drift term can be decomposed as:
@@
RFI provides an upper bound for the Wasserstein distance and the KL divergence according to HWI inequality [otto2000generalization] under the Bakry-Émery condition, with applications in information geometry and statistical physics [villani2021topics, mukherjee2018relative]. For SGLD, the RFI is given by
@@
The difference between Eq. (17) and Eq. (18) is attributed to the $\Sigma_{b}$ term, reflecting the effect of minibatch-induced stochasticity and diminishing as the batch size increases. This observation provides a theoretical justification for the empirical finding that both the sampling error and the global EPR consistently decrease with larger batch sizes. In addition, since all quantities involved ($V_{b}$, $W_{b}$, $M_{b}$, and consequently $f_{\mathrm{irr}}$ and $\Sigma_{b}$) are explicitly parameterized and estimable within our framework UIDD, Eq. (18) can be computed directly from data and thus serves as a principled scalar indicator of sampling error in SGLD.